a taxi company charges R8 up front plus R3 per kilometer travelled. If the fare is R83, what distance was travelled.
Answer by rapaljer(4671)

Let x= number of kilometers traveled
Then the total cost of the taxi is
8+3x = 83
8+3x -8 = 83-8
3x=75
x= 25 kilometers

By the way, I am from the USA where we use $. What is an R???

R^2 at SCC
Sanford, FL USA